# Joint Channel and Data Estimation for Multiuser Extremely Large-Scale MIMO Systems

### v2

@@ III Overview of the Proposed Algorithm
This section describes the overview of the proposed algorithm. The overall procedures of the proposed algorithm are illustrated in Fig. 3. As shown in the figure, the proposed algorithm mainly consists of two parts: the initial channel estimation part and subsequent JCDE part. The initial channel estimation part yields an accurate near-field channel estimate to support the convergence of the subsequent JCDE algorithm, and is composed of two stages to reduce the computational complexity. In the first stage, the angle and distance candidates from large-size polar-grids are estimated by utilizing the SOMP algorithm. In the second stage, the pairing between the path candidates obtained in the first stage and corresponding UEs is performed via the 2D-OMP algorithm by using UE specific pilot sequences. The first and second stages for initial channel estimation are described in Section IV-A and Section IV-B, respectively.
In the subsequent JCDE process, the channel and data are jointly estimated via the EP algorithm with a deterministic model-based estimation approach using the initial channel estimate. To exploit the near-field model structures, the beam-domain channel matrix $\mathbf{H}\in\mathbb{C}^{N\times U}$ is decomposed into a model-based estimate $\hat{\mathbf{S}}\in\mathbb{C}^{N\times U}$ and residual channel error $\mathbf{E}\in\mathbb{C}^{N\times U}$. $\mathbf{E}$ and $\mathbf{X}$ are jointly estimated by the EP algorithm, where the approximate joint posterior for $\mathbf{E}$ and $\mathbf{X}$ is calculated as described in Section V-C and V-D. The model-based estimate $\hat{\mathbf{S}}$ is determined by the initial channel estimate and adaptively updated in the algorithm iterations to further improve estimation performance as described in Section V-F.
Figure 3: Overflow of the proposed algorithm.
@@ IV-A Angle and Distance Estimation
To leverage the near-field channel sparsity, the virtual channel representation in the polar-domain [8] is utilized with polar-grids. The polar-grids are designed by spatially quantizing the angle and distance domain into $G_{\theta}G_{r}$ grid points as $\tilde{\bm{\theta}}\triangleq\{\tilde{\theta}_{g_{\theta}}|g_{\theta}\in\{1,\ldots,G_{\theta}\}\}$ and $\tilde{\mathbf{r}}\triangleq\left\{\tilde{r}_{g_{r},g_{\theta}}|g_{r}\in\{1,\ldots,G_{r}\},\ g_{\theta}\in\{1,\ldots,G_{\theta}\}\right\}$ with $\tilde{\theta}_{g_{\theta}}\in[-\pi,\pi]$ and $\tilde{r}_{g_{r},g_{\theta}}\in[0,\infty]$. Using the polar-grids $\tilde{\bm{\theta}}$ and $\tilde{\bm{r}}$, the polar-domain dictionary (i.e., virtual array response matrix) is designed as
$\tilde{\mathbf{A}}($ $\tilde{\bm{\theta}},\tilde{\mathbf{r}})=\left[\mathbf{a}(\tilde{\theta}_{1},\tilde{r}_{1,1}),\ldots\mathbf{a}(\tilde{\theta}_{1},\tilde{r}_{G_{r},1}),\ldots\right.$
@@
where $\tilde{\mathbf{V}}\in\mathbb{C}^{G_{\theta}G_{r}\times K_{\mathrm{p}}}$ is the row sparse matrix such that the number of nonzero rows is only $L$ and other $G_{\theta}G_{r}-L$ rows are zero since the channel is composed of a total of $L$ paths defined as in (1), with a sufficiently large number of grids, i.e., $G_{\theta}G_{r}\gg L$. Equation (8) exactly holds only if there is no quantization errors in polar grids. In actual environments, however, it approximately holds due to the presence of quantization errors. Therefore, to compensate the quantization errors, we overestimate the number of paths $\hat{L}>L$ based on the propagation environment in the considered carrier frequency [7]. To estimate $\hat{L}$ path candidates from $G_{\theta}G_{r}$ grids, the sparse reconstruction problem for $\tilde{\mathbf{V}}$ is formulated as
@@
subject $\text{subject to}\ \ \|\tilde{\mathbf{V}}\|_{2,0}=\hat{L},$ (9)
@@
The problem in (IV-A) can be approximately solved by a compressed sensing algorithm for multiple measurement vectors (MMV) problems, e.g., SOMP [8]. The computational complexity of SOMP at the $t$-th iteration in a naive implementation is $\mathcal{O}(G_{r}G_{\theta}NK_{\mathrm{p}}+Nt+Nt^{2}+t^{3})$. Its complexity can be further reduced by using the matrix inversion lemma (MIL) to $\mathcal{O}(G_{r}G_{\theta}K_{\mathrm{p}}t+Nt)$ [22, 23]. Solving the problem (IV-A) yields the angle and distance candidates corresponding to the non-zero rows of $\tilde{\mathbf{V}}$, defined as $\check{\bm{\theta}}\triangleq\{\check{\theta}_{l}\}_{l=1}^{\hat{L}}$ and $\check{\mathbf{r}}\triangleq\{\check{r}_{l}\}_{l=1}^{\hat{L}}$.
@@ IV-B UE-Path Pairing
The path candidate set $\{\check{\theta}_{l},\check{r}_{l}\}_{l=1}^{\hat{L}}$ obtained in the first stage does not specify the association of individual paths with each UE. To estimate individual channels for each UE, the second stage performs UE-path pairing, where the estimated path candidates are associated with each user using UE-specific non-orthogonal pilot sequences $\{\mathbf{x}_{\mathrm{p},u}\}_{u=1}^{U}$  [^2] In case of orthogonal piloting, one can readily imagine that this is a straightforward task.. The usage of limited path candidates $\check{\bm{\theta}}$, $\check{\mathbf{r}}$, rather than large-size polar-grids $\tilde{\bm{\theta}}$, $\tilde{\mathbf{r}}$ sampling the entire polar domain, can lead to a complexity reduction. Using the path set $\{\check{\theta}_{l},\check{r}_{l}\}_{l=1}^{\hat{L}}$, the polar-domain dictionary matrix is designed as
@@
subject $\text{subject to}\ \ \left\|\check{\mathbf{z}}\right\|_{0}=\hat{L},$ (13)
@@
The optimization problem (IV-B) is solved via a two-dimensional compressed sensing algorithm. The conventional method [10] tackles this problem with the large-size polar dictionary $\tilde{\mathbf{A}}(\tilde{\bm{\theta}},\tilde{\mathbf{r}})$ in (IV-A) instead of $\check{\mathbf{A}}(\check{\bm{\theta}},\check{\mathbf{r}})$ in (10) via the 2D-CoSaMP algorithm, which sacrifices estimation performance for complexity reduction compared to 2D-OMP [19]. In contrast, our proposed method solves the optimization problem (IV-B) via the 2D-OMP algorithm using the small-size polar-domain dictionary $\check{\mathbf{A}}(\check{\bm{\theta}},\check{\mathbf{r}})$ constructed by the path candidates $\{\check{\theta}_{l},\check{r}_{l}\}_{l=1}^{\hat{L}}$ in the first stage. As a result, the proposed method possesses the prominent capability to overcome the conventional approach [10] while retaining comparable computational complexity. Detailed discussions regarding the complexity of the proposed algorithm are presented in Section VI.
@@
5:Calculate polar-domain dictionary $\tilde{\mathbf{A}}(\tilde{\bm{\theta}},\tilde{\mathbf{r}})$ in (IV-A)
@@ V Proposed joint channel and data estimation
Given the initial estimates obtained from Algorithm 1, we aim to improve both the channel estimation performance as well as the data estimation accuracy by jointly processing the channel estimation and data detection while considering near-field properties. This section elaborates on the proposed JCDE algorithm with the initial channel estimate.
@@ V-A1 Pre-processing for Channel Estimation
To exploit the channel sparsity, the received signal and channel matrix in the spatial-domain are transformed in the beam-domain as $\mathbf{Y}\triangleq\mathbf{D}_{N}\mathbf{Y}^{\mathcal{S}}\in\mathbb{C}^{N\times K}$ and $\mathbf{H}\triangleq\mathbf{D}_{N}\mathbf{H}^{\mathcal{S}}\in\mathbb{C}^{N\times U}$ , where $\mathbf{D}_{\mathrm{N}}\in\mathbb{C}^{N\times N}$ is the DFT matrix . As described in Section II-A, the near-field channel has a cluster sparse structure due to energy leakage, thus, to tackle this problem, the channel matrix $\mathbf{H}$ is first considered as the aggregation of the model-based estimate $\hat{\mathbf{S}}\in\mathbb{C}^{N\times U}$ and the residual channel estimation error $\mathbf{E}\triangleq\mathbf{H}-\hat{\mathbf{S}}\in\mathbb{C}^{N\times U}$, resulting in
@@ V-A2 Pre-processing for Data Estimation
For low-complexity data estimation, the conventional methods based on the far-field assumption, such as [13, 12, 14], utilize MRC-based detectors that are effective in the far-field region since the beam-domain channel exhibits a simple sparse structure with a peaky spike and no correlation between the beam indices. However, these detectors are ineffective in the near-field scenario because the near-field channel has cluster sparsity due to energy leakage, and the leaked energy is correlated in the beam-domain. Although LMMSE-based detection methods such as [24, 25] are effective to deal with the correlation, these methods require matrix inversion with the size $N$, which is computationally expensive especially in XL-MIMO systems. To balance the computational complexity and detection performance, the array is virtually divided into multiple sub-arrays, and a sub-array-wise LMMSE-based detector is designed similarly to [26]. In contrast to [26], which assumes perfect channel state information (CSI), the proposed method considers the channel estimation error while jointly estimating data and channel, exploiting the near-field model structures.
@@ V-B Bayesian Inference Formulation
[…]
@@
In what follows, let us drop the iteration index $t$ for notation simplicity. The approximate posterior $g(\mathbf{E},\mathbf{X}|\mathbf{Y})$ is derived by minimizing the KL divergence subject to a Gaussian distribution set $\mathbf{\Phi}$ as
@@
These approximate factors are designed as $Q^{x}(\mathbf{X})=\prod_{c,u,k}q^{x}_{c,u,k}(x_{u,k})$, $Q^{e}(\mathbf{E})=\prod_{c}\prod_{n_{c},u,k}q^{e}_{n_{c},u,k}(e_{n_{c},u})$, $B^{x}(\mathbf{X})=\prod_{u,k}b^{x}_{u,k}(x_{u,k})$, $B^{e}(\mathbf{E})=\prod_{c}\prod_{n_{c},u}b^{e}_{n_{c},u}(e_{n_{c},u})$, where $q^{x}_{c,u,k}(\cdot)$, $q^{e}_{n,u,k}(\cdot)$, $b^{x}_{u,k}(\cdot)$, and $b^{e}_{u,k}(\cdot)$ are the parameterized approximate functions defined as
@@
where $\bm{\pi}^{q,x}_{c,u,k}\!\triangleq\!\begin{bmatrix}\hat{x}^{q}_{c,u,k},\xi^{q,x}_{c,u,k}\end{bmatrix}^{\mathrm{T}}$, $\bm{\pi}^{q,e}_{n_{c},u,k}\!\triangleq\!\begin{bmatrix}\hat{e}^{q}_{n_{c},u,k},\xi^{q,e}_{n_{c},u,k}\end{bmatrix}^{\mathrm{T}}$, $\bm{\pi}^{b,x}_{u,k}\!\triangleq\!\begin{bmatrix}\hat{x}^{b}_{u,k},\xi^{b,x}_{u,k}\end{bmatrix}^{\mathrm{T}}$, and $\bm{\pi}^{b,e}_{n_{c},u}\!\triangleq\!\begin{bmatrix}\hat{e}^{b}_{n_{c},u},\xi^{b,e}_{n_{c},u}\end{bmatrix}^{\mathrm{T}}$ are unknown parameters to be optimized by minimizing the KL divergence.
@@
Let $\mathbf{\Pi}\triangleq\{\bm{\pi}^{q,x}_{c,u,k},\ \bm{\pi}^{b,x}_{u,k},\ \bm{\pi}^{q,e}_{n_{c},u,k},\ \bm{\pi}^{b,e}_{n_{c},u}\}_{c\in\mathcal{C},n_{c}\in\mathcal{N},u\in\mathcal{U},k\in\mathcal{K}}$ denote an unknown parameter set to be optimized. The optimal unknown parameter set $\hat{\mathbf{\Pi}}$ is obtained by minimizing the KL divergence in (24). However, the objective function cannot be expressed in closed-form because $\mathrm{KL}\left(p(\mathbf{E},\mathbf{X}|\mathbf{Y};\mathbf{\Theta})\|g(\mathbf{E},\mathbf{X}|\mathbf{Y})\right)$ includes intractable integral operations with respect to the true posterior $p(\mathbf{E},\mathbf{X}|\mathbf{Y};\mathbf{\Theta})$. To tackle this, we set the target distribution $\hat{p}(\mathbf{E},\mathbf{X}|\mathbf{Y})$ instead of the true posterior $p(\mathbf{E},\mathbf{X}|\mathbf{Y};\mathbf{\Theta})$ into the KL divergence in (24). The target distribution is designed by replacing a part of the true posterior with the approximate functions in (26a)-(26d) as described in the following sections. For the sake of notation convenience for the design of the target distribution in the following section, the approximate distribution $l^{x}_{c,k}(\mathbf{E}_{c},\bar{\mathbf{x}}_{k})\simeq p(\mathbf{y}_{c,k}|\mathbf{E}_{c},\bar{\mathbf{x}}_{k})$ and $l^{e}_{n_{c},k}(\bar{\mathbf{e}}_{n_{c}},\bar{\mathbf{x}}_{k})\simeq p(y_{n_{c},k}|\bar{\mathbf{e}}_{n_{c}},\bar{\mathbf{x}}_{k})$ are expressed using (26a)-(26d) as
$l^{x}_{c,k}(\mathbf{E}_{c},\bar{\mathbf{x}}_{k})\!\propto\!\prod_{u\in\mathcal{U}}q^{x}_{c,u,k}(x_{u,k})\prod_{u\in\mathcal{U}}\prod_{n_{c}\in\mathcal{N}_{c}}\!q^{e}_{n_{c},u,k}(e_{n_{c},u}),\!\!\!$ (27a)
$l^{e}_{n_{c},k}(\bar{\mathbf{e}}_{n_{c}},\bar{\mathbf{x}}_{k})\!\propto\!\prod_{u\in\mathcal{U}}\!\!\left\{q^{x}_{c,u,k}(x_{u,k})\right\}^{\frac{1}{N_{c}}}\!\!\prod_{u\in\mathcal{U}}q^{e}_{n_{c},u,k}(e_{n_{c},u}).\!\!\!$ (27b)
To solve the KL minimization problem, the alternating optimization algorithm [30] is utilized, where a target parameter is optimized while the other parameters are fixed. In what follows, the estimation method for $\{\bm{\pi}^{q,x}_{c,u,k},\bm{\pi}^{b,x}_{u,k}\}$ and $\{\bm{\pi}^{b,e}_{n,u}$, $\bm{\pi}^{q,e}_{n,u,k}\}$ is described in Section V-C and V-D, respectively.

### v3

@@ III Overview of the Proposed Algorithm
This section describes the overview of the proposed algorithm. The overall procedures of the proposed algorithm are illustrated in Fig. 3. As shown in the figure, the proposed algorithm mainly consists of two parts: the initial channel estimation part and subsequent JCDE part. The initial channel estimation part yields an accurate channel estimate to improve the convergence point of the subsequent JCDE algorithm. This initial estimation is composed of two stages to reduce the computational complexity. In the first stage, the angle and distance candidates from large-size polar grids are estimated by utilizing the SOMP algorithm. In the second stage, the pairing between the path candidates obtained in the first stage and corresponding UEs is performed via the 2D-OMP algorithm by using UE specific pilot sequences. The first and second stages for initial channel estimation are described in Section IV-A and Section IV-B, respectively.
In the subsequent JCDE process, the channel and data are jointly estimated via the EP algorithm with a deterministic model-based estimation approach, employing the initial channel estimate as the starting values. To exploit the near-field model structures, the beam-domain channel matrix $\mathbf{H}\in\mathbb{C}^{N\times U}$ is decomposed into a model-based estimate $\hat{\mathbf{S}}\in\mathbb{C}^{N\times U}$ and residual channel error $\mathbf{E}\in\mathbb{C}^{N\times U}$. $\mathbf{E}$ and $\mathbf{X}$ are jointly estimated by the EP algorithm, where the approximate joint posterior for $\mathbf{E}$ and $\mathbf{X}$ is calculated as described in Section V-C and V-D. The model-based estimate $\hat{\mathbf{S}}$ is determined by the initial channel estimate and adaptively updated in the algorithm iterations to further improve estimation performance as described in Section V-F. The performance of data and channel estimation in the JCDE algorithm heavily depends on the accuracy of the initial channel estimate. Therefore, creating an accurate initial channel estimate leads to a better convergence point in the JCDE algorithm, which is the main role of the initial channel estimation algorithm.
Figure 3: Overview of the proposed algorithm.
@@ IV-A Angle and Distance Estimation
To leverage the near-field channel sparsity, the virtual channel representation in the polar-domain [7] is utilized with polar grids. The polar grids are designed by spatially quantizing the angle and distance domain into $G_{\theta}G_{r}$ grid points as $\tilde{\bm{\theta}}\triangleq\{\tilde{\theta}_{g_{\theta}}|g_{\theta}\in\{1,\ldots,G_{\theta}\}\}$ and $\tilde{\mathbf{r}}\triangleq\left\{\tilde{r}_{g_{r},g_{\theta}}|g_{r}\in\{1,\ldots,G_{r}\},\ g_{\theta}\in\{1,\ldots,G_{\theta}\}\right\}$ with $\tilde{\theta}_{g_{\theta}}\in[-\pi/2,\pi/2]$ and $\tilde{r}_{g_{r},g_{\theta}}\in[0,\tilde{r}_{\mathrm{max}}]$333 In this paper, we design the polar grids using the near-field sampling method proposed in [9]. Using this sampling method, the maximum distance grid is computed as $\tilde{r}_{\mathrm{max}}=2N^{2}\lambda\sqrt{\frac{0.001624}{1-\bar{\gamma}_{\mathrm{coh}}}}$ with a desired coherence value $\bar{\gamma}_{\mathrm{coh}}$, which is set to 0.6 in the numerical simulation in Section VI. . Using the polar grids $\tilde{\bm{\theta}}$ and $\tilde{\bm{r}}$, the polar-domain dictionary (i.e., virtual array response matrix) is designed as
[…]
@@
where $\tilde{\mathbf{V}}\in\mathbb{C}^{G_{\theta}G_{r}\times K_{\mathrm{p}}}$ is the row sparse matrix such that the number of nonzero rows is only $L$ and other $G_{\theta}G_{r}-L$ rows are zero since the channel is composed of a total of $L$ paths defined as in (1), with a sufficiently large number of grids, i.e., $G_{\theta}G_{r}\gg L$. Equation (8) exactly holds only if there is no quantization errors in polar grids. In actual environments, however, it approximately holds due to the presence of quantization errors. Therefore, to compensate the quantization errors, we overestimate the number of paths $(\hat{L}>L)$444 This paper assumes that prior knowledge of the number of paths can be obtained from long-term statistics because the temporal variation in the number of paths is significantly slower than that of the channel. This assumption has been widely adopted in various studies [6, 22]. . To estimate $\hat{L}$ path candidates from $G_{\theta}G_{r}$ grids, the sparse reconstruction problem for $\tilde{\mathbf{V}}$ is formulated as
@@
$\text{subject to}\ \ \|\tilde{\mathbf{V}}\|_{2,0}=\hat{L}$, (9)
@@
The problem in (IV-A) can be approximately solved by a compressed sensing algorithm for multiple measurement vectors (MMV) problems, e.g., SOMP [7], because $\tilde{\mathbf{V}}$ has common support across its columns (i.e., along the pilot dimension), meaning that the index set of nonzero elements is identical for all column vectors of $\tilde{\mathbf{V}}$. The complexity of SOMP at the $t$-th iteration in a naive implementation is $\mathcal{O}(G_{r}G_{\theta}NK_{\mathrm{p}}+Nt+Nt^{2}+t^{3})$. Its complexity can be further reduced by using the matrix inversion lemma (MIL) to $\mathcal{O}(G_{r}G_{\theta}K_{\mathrm{p}}t+Nt)$ [23, 24]. Solving the problem (IV-A) yields the angle and distance candidates corresponding to the non-zero rows of $\tilde{\mathbf{V}}$, defined as $\check{\bm{\theta}}\triangleq\{\check{\theta}_{l}\}_{l=1}^{\hat{L}}$ and $\check{\mathbf{r}}\triangleq\{\check{r}_{l}\}_{l=1}^{\hat{L}}$.
@@ IV-B UE-Path Pairing
The path candidate set $\{\check{\theta}_{l},\check{r}_{l}\}_{l=1}^{\hat{L}}$ obtained in the first stage does not specify the association of individual paths with each UE. To estimate individual channels for each UE, the second stage performs UE-path pairing, where the estimated path candidates are associated with each user using UE-specific non-orthogonal pilot sequences $\{\mathbf{x}_{\mathrm{p},u}\}_{u=1}^{U}$. The usage of limited path candidates $\check{\bm{\theta}}$, $\check{\mathbf{r}}$, rather than large-size polar grids $\tilde{\bm{\theta}}$, $\tilde{\mathbf{r}}$ sampling the entire polar-domain, can lead to a complexity reduction. Using the path set $\{\check{\theta}_{l},\check{r}_{l}\}_{l=1}^{\hat{L}}$, the polar-domain dictionary matrix is designed as
@@
$\text{subject to}\ \ \left\|\check{\mathbf{z}}\right\|_{0}=\hat{L}$, (13)
@@
[…]
@@
[…]
@@ V Proposed joint channel and data estimation
Given the initial estimates obtained from Algorithm 1, we aim to improve both the channel and data estimation performance while considering near-field properties. This section elaborates on the proposed JCDE algorithm using the initial channel estimate as the starting values.
@@ V-A1 Pre-processing for Channel Estimation
To exploit the channel sparsity, the received signal and channel matrix in the spatial-domain are transformed in the beam-domain as $\mathbf{Y}\triangleq\mathbf{D}_{N}\mathbf{Y}^{\mathcal{S}}\in\mathbb{C}^{N\times K}$ and $\mathbf{H}\triangleq\mathbf{D}_{N}\mathbf{H}^{\mathcal{S}}\in\mathbb{C}^{N\times U}$. Since the DFT matrix $\mathbf{D}_{N}$ is unitary, the transformation between the spatial and beam-domains can be efficiently performed via simple matrix multiplication unlike the polar-domain representation. Therefore, the matrix size remains unchanged after the transformation, resulting in no additional computational complexity in the subsequent iterative channel estimation and data detection processes. However, as described in Section II-A, the near-field channel in the beam-domain has a cluster sparse structure due to energy leakage. Thus, to tackle this problem, the channel matrix $\mathbf{H}$ is first considered as the aggregation of the model-based estimate $\hat{\mathbf{S}}\in\mathbb{C}^{N\times U}$ and the residual channel estimation error $\mathbf{E}\triangleq\mathbf{H}-\hat{\mathbf{S}}\in\mathbb{C}^{N\times U}$, resulting in
@@ V-A2 Pre-processing for Data Estimation
For low-complexity data estimation, the conventional methods based on the far-field assumption, such as [12, 11, 13], utilize MRC-based detectors, which are ineffective in the near-field scenario because the near-field channel has cluster sparsity due to energy leakage, and the leaked energy is correlated in the beam-domain. Although LMMSE-based detection methods such as [25, 26] are effective to deal with the correlation, these methods require matrix inversion with the size $N$, which is computationally expensive especially in XL-MIMO systems. To balance the computational complexity and detection performance, the array is virtually divided into multiple sub-arrays, and a sub-array-wise LMMSE-based detector is designed similarly to [27]. In contrast to [27], which assumes perfect channel state information (CSI), the proposed method considers the channel estimation error while jointly estimating data and channel, exploiting the near-field model structures.
@@ V-B Bayesian Inference Formulation
$\text{E-step : }\mathcal{F}(\mathbf{\Theta},\mathbf{\Theta}^{(t)})\!=\!\mathbb{E}_{p(\mathbf{E},\mathbf{X}|\mathbf{Y};\mathbf{\Theta}^{(t)})}\!\left[\ln p(\mathbf{Y},\mathbf{E},\mathbf{X};\mathbf{\Theta})\right]+\mathsf{c}_{0}^{(t)}\!,$
$\text{M-step : }\mathbf{\Theta}^{(t+1)}=\underset{\mathbf{\Theta}}{\mathrm{argmax}}\ \mathcal{F}(\mathbf{\Theta},\mathbf{\Theta}^{(t)}),$ (22)
@@
For notation simplicity, let us drop the iteration index $t$. The approximate posterior $g(\mathbf{E},\mathbf{X}|\mathbf{Y})$ is derived by KL minimization subject to a Gaussian distribution set $\mathbf{\Phi}$ as
@@
These approximate factors are designed as $Q^{x}(\mathbf{X})=\prod_{c,u,k}q^{x}_{c,u,k}(x_{u,k})$, $Q^{e}(\mathbf{E})=\prod_{c}\prod_{n_{c},u,k}q^{e}_{n_{c},u,k}(e_{n_{c},u})$, $B^{x}(\mathbf{X})=\prod_{u,k}b^{x}_{u,k}(x_{u,k})$, $B^{e}(\mathbf{E})=\prod_{c}\prod_{n_{c},u}b^{e}_{n_{c},u}(e_{n_{c},u})$, where $q^{x}_{c,u,k}(\cdot)$, $q^{e}_{n_{c},u,k}(\cdot)$, $b^{x}_{u,k}(\cdot)$, and $b^{e}_{u,k}(\cdot)$ are the parameterized approximate functions defined as
@@
where $\bm{\pi}^{q,x}_{c,u,k}\!\triangleq\!\begin{bmatrix}\hat{x}^{q}_{c,u,k},\xi^{q,x}_{c,u,k}\end{bmatrix}^{\mathrm{T}}$, $\bm{\pi}^{q,e}_{n_{c},u,k}\!\triangleq\!\begin{bmatrix}\hat{e}^{q}_{n_{c},u,k},\xi^{q,e}_{n_{c},u,k}\end{bmatrix}^{\mathrm{T}}$, $\bm{\pi}^{b,x}_{u,k}\!\triangleq\!\begin{bmatrix}\hat{x}^{b}_{u,k},\xi^{b,x}_{u,k}\end{bmatrix}^{\mathrm{T}}$, and $\bm{\pi}^{b,e}_{n_{c},u}\!\triangleq\!\begin{bmatrix}\hat{e}^{b}_{n_{c},u},\xi^{b,e}_{n_{c},u}\end{bmatrix}^{\mathrm{T}}$ are unknown parameters to be optimized by KL minimization.
@@
Let $\mathbf{\Pi}\triangleq\{\bm{\pi}^{q,x}_{c,u,k},\ \bm{\pi}^{b,x}_{u,k},\ \bm{\pi}^{q,e}_{n_{c},u,k},\ \bm{\pi}^{b,e}_{n_{c},u}\}_{c\in\mathcal{C},n_{c}\in\mathcal{N},u\in\mathcal{U},k\in\mathcal{K}}$ denote an unknown parameter set. The optimal parameter set $\hat{\mathbf{\Pi}}$ is obtained by KL minimization in (23). However, the objective function cannot be expressed in closed-form because $\mathrm{KL}\left(p(\mathbf{E},\mathbf{X}|\mathbf{Y};\mathbf{\Theta})\|g(\mathbf{E},\mathbf{X}|\mathbf{Y})\right)$ includes intractable integral with respect to the true posterior $p(\mathbf{E},\mathbf{X}|\mathbf{Y};\mathbf{\Theta})$. To tackle this, we set the target distribution $\hat{p}(\mathbf{E},\mathbf{X}|\mathbf{Y})$ instead of the true posterior into the KL divergence in (23). The target distribution is designed by replacing a part of the true posterior with the approximate functions in (26a)-(26d) as described in the following sections. For the sake of notation for the design of the target distribution, the approximate distribution $l^{x}_{c,k}(\mathbf{E}_{c},\bar{\mathbf{x}}_{k})\simeq p(\mathbf{y}_{c,k}|\mathbf{E}_{c},\bar{\mathbf{x}}_{k})$ and $l^{e}_{n_{c},k}(\bar{\mathbf{e}}_{n_{c}},\bar{\mathbf{x}}_{k})\simeq p(y_{n_{c},k}|\bar{\mathbf{e}}_{n_{c}},\bar{\mathbf{x}}_{k})$ are expressed using (26a)-(26d) as
$\hskip-6.45831ptl^{x}_{c,k}(\mathbf{E}_{c},\bar{\mathbf{x}}_{k})\!\propto\!\prod_{u\in\mathcal{U}}q^{x}_{c,u,k}(x_{u,k})\prod_{u\in\mathcal{U}}\prod_{n_{c}\in\mathcal{N}_{c}}\!q^{e}_{n_{c},u,k}(e_{n_{c},u}),\!\!\!$ (27a)
$\hskip-6.45831ptl^{e}_{n_{c},k}(\bar{\mathbf{e}}_{n_{c}},\bar{\mathbf{x}}_{k})\!\propto\!\prod_{u\in\mathcal{U}}\!\!\left\{q^{x}_{c,u,k}(x_{u,k})\right\}^{\frac{1}{N_{c}}}\!\!\prod_{u\in\mathcal{U}}q^{e}_{n_{c},u,k}(e_{n_{c},u}).\!\!\!$ (27b)
To solve the KL minimization problem, the alternating optimization algorithm [30] is utilized, where a target parameter is optimized while the other parameters are fixed. In what follows, the estimation methods for $\{\bm{\pi}^{q,x}_{c,u,k},\bm{\pi}^{b,x}_{u,k}\}$ and $\{\bm{\pi}^{b,e}_{n,u}$, $\bm{\pi}^{q,e}_{n,u,k}\}$ are described in Section V-C and V-D, respectively.